OpenStudy (anonymous):

Solve. Express the answer in simplest form. - 1/3 + w = 5 /38

OpenStudy (anonymous):

@science00000

OpenStudy (anonymous):

5 3/8 I don't know how to edit

OpenStudy (anonymous):

Like last time, you want to isolate the variable.

OpenStudy (anonymous):

\[w = 5\frac{ 3 }{ 8 }+\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

Make both fractions into improper form and common denominator. \[w=\frac{ 43 }{ 8 }+\frac{ 1 }{ 3 }\] Common denominators \[w = \frac{ 129 }{ 24} + \frac{ 8 }{ 24 }\] Which would equal to \[w = \frac{ 137 }{ 24 }\]

OpenStudy (anonymous):

Then you Simply?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

It's hard to find a pencil in the dark

OpenStudy (anonymous):

\[w=5\frac{ 17 }{ 24 }\]

OpenStudy (anonymous):

I got three more then I might post more

OpenStudy (anonymous):

I have to go now :)

OpenStudy (anonymous):

AWwwe